OpenStudy (anonymous):

Can someone please help me solve this problem 7/(x^2+10x+21)+2/(x+3)=1/(x^2+10x+21) I'd also like to know how to solve it as well. This is so confusing, thanks

OpenStudy (campbell_st):

can you check the question as there appears to be something wrong here is what I would do well the factor the quadratic denominator \[\frac{7}{(x +3)(x + 7)} + \frac{2}{x + 3}\] you need a common denominator so multiply the 2nd fraction by (x + 7)/(x+7) and you get \[\frac{7}{(x +3)(x+7)} + \frac{2}{(x+3)} \times \frac{(x + 7)}{(x + 7)} = \frac{7 + 2(x + 7)}{(x +3)(x+7)}\] then simplify from here..

OpenStudy (anonymous):

I want to help, but really don't have time. : ( http://www.wolframalpha.com/ This website will help. Just enter the formula in, and it will give you the answer! Sorry, hope this helps!